OpenStudy (anonymous):

simplify the expression (x+6)^2

OpenStudy (anonymous):

SO WHAT'S THE FIST STEP??

OpenStudy (anonymous):

Idk haven't did this in a while

OpenStudy (yoloshroom):

Alright so, ^2 mean's it's multiplied by itself. so (x+6)^2 = (x+6)(x+6) Can you solve from there?

OpenStudy (theedwardsfamily):

What do you think?

OpenStudy (anonymous):

x2+12x+36

OpenStudy (theedwardsfamily):

You can try this: https://www.symbolab.com/

OpenStudy (theedwardsfamily):

@Ymcagymnasticas123

OpenStudy (anonymous):

WRONG OR RIGHT?

OpenStudy (yoloshroom):

Symbolab <3 in college and still using it.. www.mathway.com i prefer that, but it doesnt give free steps

OpenStudy (theedwardsfamily):

Yea, ikr! lol

OpenStudy (anonymous):



OpenStudy (anonymous):

thanks

OpenStudy (mathstudent55):

You can think of this problem two ways. 1. Squaring a binomial is simply multiplying the binomial by itself. \((x + 6)^2 = (x + 6)(x + 6)\) Now use FOIL, for example, to find the product. 2. You can memorize the pattern of squaring a binomial: \((a + b)^2 = a^2 + 2ab + b^2\) Now just follow the pattern and find the answer.